Blauser's started the year with $280 in cash, $924 in inventory, $361 in accounts payable, $1,687 in equipment, and $414 in accounts receivable. At year's end, the firm had $311 in cash, $1,594 in equipment, $1,003 in inventory, $426 in accounts receivable, and $398 in accounts payable. What was the change in net working capital during the year?

A. -$860

B $191

C $85

D -$94

E -$206
Answer #1

Blauser's

Working capital = Current assets - Current liabilities

Current assets = Cash + Inventory + Accounts receivable ...etc

Current liabilities = Accounts payable ...etc

Opening working capital = (280+924+414) - 361

Opening working capital = $ 1,257

Closing working capital = (311+1,003+426) - 398

Closing working capital = $ 1,342

Net Change in working capital during the year = Closing working capital - Opening working capital

Net change in working capital = 1,342 - 1,257 = $ 85

Hence, Option "C" is correct.